OpenStudy (anonymous):

Complete the square to solve for x. x^2−4x−32=0

OpenStudy (anonymous):

My answer is (x-2)^2=36 x=38 or x=-34 is it correct?

OpenStudy (shamil98):

(x-2)^2 = 36 (x-2) = 6 x = 8 through completing the square.

OpenStudy (shamil98):

How did you end up with x = 38 and x = -34? ..